How many solutions are in -3x+8=2x+9-5x
Answer by Edwin McCravy(20054)

 -3x+8 = 2x+9-5x

Combine the 2x and the -5x on the right side and get -3x

 -3x+8 = -3x+9

Add 3x to both sides

 -3x+8 = -3x+9
 +3x     +3x
---------------
  0x+8 =  0x+9

Since 0x is just 0, we have

     8 =     9

Since 8 does not equal to 9, the equation is false and there
can be no solutions.

So there are 0 solutions.

Edwin